OpenStudy (anonymous):

Choose the correct simplification of the expression (8x2)3. a)24x6 b)512x6 c)512x5 d)24x5

jhonyy9 (jhonyy9):

do you know the answer on this (a^2)^3 = ?

jhonyy9 (jhonyy9):

so this is equal (a^2)^3 =a^(2*3) =a^6 ok ?

jhonyy9 (jhonyy9):

so than in case of your exercise ?

jhonyy9 (jhonyy9):

will be 8^3 time (x^2)^3 = can you calcule it ?

jhonyy9 (jhonyy9):

8^3 =8*8*8=64*8=512 and (x^2)^3 =x^(2*3)=x^6 so than the right answer will be ?

jhonyy9 (jhonyy9):

thank you saifoo.khan there are ,,students" anytime who dont like,wann cooperating nothing

jhonyy9 (jhonyy9):

just getting the right answer